3. The following table gives rounded annual GDP figures as of the third quarter (July-Aug-Sep)

2018:

Consumption spending

$14,050

Investment spending

$3,710

Depreciation

$1,750

Government spending on goods and services

$3,550

Government transfer payments

$8,200

Total exports

$2,540

Total imports

$3,195

a. Calculate GDP using the spending approach. (Note that not all values will be used.)

b. Calculate the percentage of each spending component relative to total GDP. Are your percentages consistent with what we should expect for the US economy? Explain.
Answer #1

a) GDP Expenditure approach= C+I+G+NX

Consumption Spending+Investment Spending+Government spending on goods and services+(Exports-Imports)

=14, 050+3710+3550+ ( 3,195-2540)

= 21965

b) Percentage of each spending component

Consumption spending=14050*100/21965=63.96%

Investment Spending = 3710*100/21965= 16.89%

Government Spending on goods =3550*100/21965=16.16%

Total Exports = 2540*100/21965=11.56%

Total Import = 3195*100/21965 = 14.54%

No all our percentage are not consistent they differ from each other, since percentage consumption was more than investment and imports are more than exports. In order to improve the condition economy needs to focus more on Exports rather than Imports.